OpenStudy (anonymous):

tell how many solutions the system of equations would have 6x - 2y = -8

OpenStudy (pinksapphire):

All solutions.

OpenStudy (yanasidlinskiy):

Do you know how to solve this equation?

OpenStudy (pinksapphire):

6x-2y=-8 -6x -6x Subtract 6 from both sides. -2y/-2=-8-6x/-2 Divide by -2. y=4+3x Now you have y. 6x-2(4+3x)=-8 Plug in 4+3x into y in 6x-2y=-8. 6x-8-6x=-8 Since -8 + 8 = 0 cancel the number out. +8 +8 6x-6x=0 Since 6x-6x it will become 0 so you just cancel it out. 0=0 Every time you get 0=0, this means that any number can be a solution.

OpenStudy (yanasidlinskiy):

Nice work!